OpenStudy (anonymous):

Sam bought brushes for $8 , a palette for $5 and oil paints for $15 he paid $29.82 in all what sales tax rate did Sam pay

OpenStudy (anonymous):

Sam's total purchase is 8 + 5 + 15 = 28 He paid 29.82. So, algebraically, 28 * (1+ r) = 29.82 28 + 28r = 29.82 28r = 1.82 r = 1.82/28 r = .065 or 6.5% sales tax. Prove it: 28 * (1 + .065) = 29.82?? 28 x 1.065 = 29.82?? 29.82 = 29.82 YES!! Hope that helps.